OpenStudy (anonymous):

Find the vertex of the function. Write your answer in the form (x,y). Use the slash mark ( / ) as a fraction bar if necessary but do not use decimal points. x^2+6x+7

OpenStudy (across):

x^2+6x+7=(x+7)(x-1) (1-7)/2=-3 the vertex is at (-3,-2)

OpenStudy (anonymous):

thank yoU